f (-3,2) lies on the circle x2+y2+2gx+2fy+c=0 which is concentric with the circle x2+y2+6x+8y-5=0 then c=?

Grade:12

1 Answers

Arun
5 years ago
Dear Suhas
 
 
Circle will be-
 
x² +y² + 6x + 8y jc = 0
 
If (-3,2) lies on this circle
 
9 + 4 -18 +16 +c = 0
 
c +11 = 0
 
c = -11
 
 
Regards
Arun (askIITians forum expert)